For the set of data below, determine the 5 number summaries. Construct the resulting box and whisker plot.
5, 3, 8, 6, 1, -10, 3, 9, 7, 4, 2,
1
Expert's answer
2020-10-19T18:15:44-0400

 Put the values in the data set into increasing order: 


"-10,1,2,3,3,4,5,6,7,8,9"

The minimum (the smallest number) is "-10" and the maximum (the largest number) is "9."

Find the median. The median is the middle number. The median is "4."

Find Q1 and Q3.


"Q1=2, Q3=7"

Minimum"=-10", "Q1=2", median "=4", "Q3=7", maximum "=9".

"Range=max-min=9-(-1)=19"

"Interquartile \ range = Q3 -Q1=7-2=5."